OpenStudy (anonymous):

Okey so I'm suppose to do these equation with substitution and can you guys lakes help me! 6x – 9y = 12 x + y = -18

OpenStudy (mathstudent55):

To solve a system, of equations using the substitution method, you must first solve one equation for one variable.

OpenStudy (zzr0ck3r):

solve for y in the second equation \(y=-18-x\) now plug that into the first equation for y \(6x-9(-18-x)=12\) solve for x

OpenStudy (mathstudent55):

Here are the two equations of the system of equations: 6x – 9y = 12 x + y = -18

OpenStudy (mathstudent55):

Before you even use the substitution method, you can simplify the first equation by dividing both sides by 3 since every coefficient is divisible by 3. Then the system of equations becomes: 2x – 3y = 4 x + y = -18

OpenStudy (mathstudent55):

Now you need to solve one equation for one variable. You can pick either equation and either variable, but it's faster if you pick an equation and a variable that are easy to solve for.

OpenStudy (mathstudent55):

Looking at the second equation, you see that the variables x and y don't have coefficients (numbers multiplying them), so pick either x or y and the second equation.

OpenStudy (mathstudent55):

Let's say we choose to solve the second equation for x. We need to isolate x, so we subtract y from both sides. 2x – 3y = 4 ----(stays the same)----> 2x - 3y = 4 x + y = -18 ----(solved for x)-------> x = -18 - y

OpenStudy (mathstudent55):

Now we do the substitution step. We take the first equation, and we substitute what x is equal to (from the second equation) into the first equation. 2(-18 - y) - 3y = 4 Now we have one equation with only one variable (y), so we can solve for y. -36 - 2y - 3y = 4 -36 - 5y = 4 -5y = 40 y = -8 Now we plug in this value of y into either of the original equations and solve for x. I'll use the second equation because it's easier. x + y = -18 x + (-8) = -18 x - 8 = -18 x = -10 The solution is: x = -10, y = -8

OpenStudy (mathstudent55):

When you read this, read it a few times and study it. This is how the substitution method works.

OpenStudy (mathstudent55):

If you have any questions, post them. I'll try to answer when I see them.